#37 What are the correct yields in relation to the Treasury bill described below?

Tenor: 180 days

Face value: $185,000

Selling price: $167,250

A. Bank Discount Yield Money market Yield Effective Annual Yield

10.61% 12.36% 12.83%

B. Bank Discount Yield Money Market Yield Effective Annual Yield

19.19% 21.52%

C. Bank Discount Yield Money Market Yield Effective Annual Yield

15.72% 17.08% 17.69%
Answer #1

Bank Discount yield = (F-P)*360/(F*T) ,where F = face value , P = selling price , T= tenor days

= (185,000- 167250)*360/(185,000*180) = 19.19%

Money Market Yield = (360* BDY)/ (360- T*BDY) , where BDY is bank discount yield

= (360*0.1919)/ (360- 180*0.1919)

= 21.52%

hence option B is the right answer